OpenStudy (anonymous):

I'm working on my homework on 3D vectors and the instruction says Determine (ordered triple) vector |AB| and its magnitude |AB| for given points: A(2,5,4) B(-1,1,2) Can someone help please?!

OpenStudy (oaktree):

So the vector itself is just the distance change for each x, y, and z component. So the vector is\[<(2-(-1)),(5-1),(4-2)> = <3,4,2>\]Now that we have the vector, the magnitude is easy to calculate.\[|<3,4,2>| = \sqrt{3^2 + 4^2 + 2^2} = \sqrt{9+16+4} = \sqrt{29}\]Makes sense?

OpenStudy (anonymous):

Oh my gosh! Yes!! Thank you so much!!